OpenStudy (cookieeater123):

please help An equation is shown below: 5(2x - 8) + 15 = -15 Write the steps you will use to solve the equation and explain each step.

OpenStudy (wanell):

5(2x-8)+15=-15 One solution was found : x = 1 Rearrange: Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : 5*(2*x-8)+15-(-15)=0 Step by step solution : Step 1 : Step 2 : Pulling out like terms : 2.1 Pull out like factors : 2x - 8 = 2 • (x - 4) Equation at the end of step 2 : (10 • (x - 4) + 15) - -15 = 0 Step 3 : Step 4 : Pulling out like terms : 4.1 Pull out like factors : 10x - 10 = 10 • (x - 1) Equation at the end of step 4 : 10 • (x - 1) = 0 Step 5 : Equations which are never true : 5.1 Solve : 10 = 0 This equation has no solution. A a non-zero constant never equals zero. Solving a Single Variable Equation : 5.2 Solve : x-1 = 0 Add 1 to both sides of the equation : x = 1 One solution was found : x = 1 Processing ends successfull

OpenStudy (didyouseethat):

5(2x + -8) + 15 = -15 5(-8 + 2x) + 15 = -15 (-8 * 5 + 2x * 5) + 15 = -15 (-40 + 10x) + 15=-15 Reorder the terms: -40 + 15 + 10x = -15 Combine like terms: -40 + 15 = -25 -25 + 10x = -15 -25 + 10x = -15 Solve for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '25' to each side of the equation. -25 + 25 + 10x = -15 + 25 Combine like terms: -25 + 25 = 0 0 + 10x = -15 + 25 10x = -15 + 25 Combine like terms: -15 + 25 = 10 10x = 10 Divide each side by '10'. x = 1 Simplify x = 1

OpenStudy (cookieeater123):

thank's for the help

OpenStudy (wanell):

NP